OpenStudy (anonymous):

Areas between graphs of 2 functions

OpenStudy (anonymous):

y=x^2, x+y=2

OpenStudy (anonymous):

Alright, I'll lead you through the steps, how about that? :)

OpenStudy (anonymous):

1) Sketch a graph of both functions

OpenStudy (anonymous):

sounds fantastic! thank you! lol

OpenStudy (anonymous):

np ^_^

OpenStudy (anonymous):

graph was given lol (way easier)

OpenStudy (anonymous):

lol, okay!

OpenStudy (anonymous):

2) find the intersection points when y1= y2

OpenStudy (anonymous):

set equal to zero?

OpenStudy (anonymous):

yep, find the x's :)

OpenStudy (anonymous):

got em...-2 and 1

OpenStudy (anonymous):

alright, good :)

OpenStudy (anonymous):

now, we must find out which function is bigger :) in other words, which graph is on the top of the other?

OpenStudy (anonymous):

I think y= x^2 is on top of the other one, right? :)

OpenStudy (anonymous):

yeah i remember that from class but it sort of looks like an upside down "V" so neither is on top. Do i use the right one because its a graph of y not x?

OpenStudy (anonymous):

exactly! you can see the intersection in the graph right? which curve is on the top?

OpenStudy (anonymous):

i think for y he said right-left and for x top-bottom

OpenStudy (anonymous):

let me sketch the graph lol just to make sure :)

OpenStudy (anonymous):

yep i see the intersection lol thanks. i would upload the pic but that would take longer than a sketch

OpenStudy (anonymous):

hold on :)

OpenStudy (anonymous):

okay got it, look at your sketch, color the region bounded (where the intersection lies/ closed region lol)

OpenStudy (anonymous):

which line is on the top? :)

OpenStudy (anonymous):

in my sketch, I see y = 2-x on top :)

OpenStudy (anonymous):

hmm, I got a different sketch , I've sovled for y for the second equation (y=2-x)

OpenStudy (anonymous):

try to sketch the graph using y = x^2 and y = 2-x :)

OpenStudy (anonymous):

A plot related to this problem is attached.

OpenStudy (anonymous):

yeah, that's what I got robto ^_^, birdsell are you following?

OpenStudy (anonymous):

its the same graph its just the given one is a snippet

OpenStudy (anonymous):

LOL, then draw your own to make your life easier :) , look at robto's sketch , which line is on the top ? ^_^

OpenStudy (anonymous):

intersection point of the 2 lines in Q1 is at 1,1

OpenStudy (anonymous):

are you sure? I got x = -2 and x = 1

OpenStudy (anonymous):

so did i :-X but the picture shows it at 1,1

OpenStudy (anonymous):

birdsell, dear :) , let's take it step by step 1) sketch both graphs, on your own , ignore the given :)

OpenStudy (anonymous):

you have y =x^2 and y = 2-x

OpenStudy (anonymous):

sketch them

OpenStudy (anonymous):

you'll have to get the same sketch as robto's ^_^

OpenStudy (anonymous):

im so confused :( most of them are clear that one is on top of the other

OpenStudy (anonymous):

calm down, did you draw the graph?

OpenStudy (anonymous):

and im attempting to use my calculator to graph the points but cannnt

OpenStudy (anonymous):

don't! you can try out values, alright for example, sketch y = x^2 first, take values for x and find y, then plot them :) same with the other equation, do that

OpenStudy (anonymous):

after that, you'll get a graph similar to robto's

OpenStudy (anonymous):

try :)

OpenStudy (anonymous):

The second plot version uses the same formulas except the vertical and horizontal scaling is the same.

OpenStudy (anonymous):

so x^2 is on top?

OpenStudy (anonymous):

nope, look at the sketch again :)

OpenStudy (anonymous):

The curve passing through (0,0) is x^2.

OpenStudy (anonymous):

look at the closed region,=)

OpenStudy (anonymous):

lmfao im dumb :)... got it

OpenStudy (anonymous):

no you're not, you panicked :D, alright so which line is on the top?

OpenStudy (anonymous):

y=2-x

OpenStudy (anonymous):

excellent, so this wil be you g(x) :) and the other curve is f(x) okay?

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

I actually think I got it! lol wanna help with another?

OpenStudy (anonymous):

now to find the Area, the general formula is :\[A =\int\limits_{a}^{b}[g(x) - f(x)] dx\] where g(x) is the curve that's in the top and f(x) in the bottom, replace the equations and solve : \[A = \int\limits_{-2}^{1} (2-x) -(x^2) dx\] \[A = \int\limits_{-2}^{1}( 2-x-x^2)dx\] \[A = [2x-x^2/2 + x^3/3]\] from -2 to 1 so you'll get : \[A = [2(1)-(1)^2/2-(1)^3/3] - [2(-2) -(-2)^2/2 - (-2)^3/3]\] compute it and you'll get A lol :)

OpenStudy (anonymous):

sure , I'll be glad to help ^_^

OpenStudy (anonymous):

hanks! new thread or current?

OpenStudy (anonymous):

thanks**

OpenStudy (anonymous):

lol whatever you want :), np ^_^

OpenStudy (anonymous):

not sure if you are familiar with slicing?

OpenStudy (anonymous):

yes I am , it has to do with finding the volume right =)?

OpenStudy (anonymous):

yep! :):)

OpenStudy (anonymous):

I had a midterm on it last sunday lol!

OpenStudy (anonymous):

alright, which method does it want you to use? :)

OpenStudy (anonymous):

ughhh college... calc is killin me... up until now i have been great

OpenStudy (anonymous):

lol, calculus II right? practice, work hard, you'll perfect everything. Work hard now to relax in the end, believe me I sucked in this part =-= but now it's just SIMPLE :)

OpenStudy (anonymous):

Sorry to intterrupt, your question ask area of between x^2 , y+x=2 and x axis according to photo but the solution answers area of two function.I am confused

OpenStudy (anonymous):

you're finding the area of the 2 functions between the 2 curves Hakan, to do that you have to find the enclosed region :)

OpenStudy (anonymous):

actually calc 1. I go to a tough school and this is calc 1 carriculum

OpenStudy (anonymous):

serious? I'm taking this in calculus II in college lol

OpenStudy (anonymous):

Ill start a new thread sstarcia

OpenStudy (anonymous):

well, I took the area b/w the curves in calculus I but the volume in calculus II

OpenStudy (anonymous):

alright :)

OpenStudy (anonymous):

ahhh it isnt working lol

OpenStudy (anonymous):

how come?

OpenStudy (anonymous):

anyway... The solid lies between planes perpendicular to the y axis at y=0 and y=2 the cross-sections perpendicular to the y-axis are circular discs with diameters running from the y-axis to the parabola \[x=\sqrt{5y^2}\]

OpenStudy (anonymous):

But the photo of question doesn't ask that as i see.Anyway.Thanx :)

OpenStudy (anonymous):

lol we re-sketched it, look at robto's sketch :) np :)

OpenStudy (anonymous):

alright, to make your life easier bird, find y of this equation

OpenStudy (anonymous):

\[y = x/\sqrt(5)\] I guess

OpenStudy (anonymous):

i got y=x/5 lol

OpenStudy (anonymous):

lol, take y^2 out and you'll get y, then divide by sqrt of 5 :)

OpenStudy (anonymous):

nvm i see now lol gotta take Sqrt of both top and bottom

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

now: 1) sketch the given in a graph, y=0, y=2 and y=x/sqrt(5)

OpenStudy (anonymous):

use your calculator for y = x/sqrt(5) you'll get something like a slight curve on the right, right?

OpenStudy (anonymous):

it wont give me anything... when i type y=x/sqrt[5]

OpenStudy (anonymous):

lol, forget your graphying calculator and do it by hand. Take values for x and calculate y using the calculator :)

OpenStudy (anonymous):

graphing*

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

I think you'll get something like this , I think

OpenStudy (anonymous):

did you get something similar to it?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

alright, now what did the question say, how does the curve rotate? about the y-axis, x-axis?